OpenStudy (anonymous):

Solve for x: (x - 9)2 = 1 by completing the square

OpenStudy (radar):

\[(x-9)^{2}=1\]It is already complete, only need to solve: Taking square root of both sides getting:\[x-9=\pm1\]\[x-9=1\]\[x=10\]\[x-9=-1\]\[x=-8\]